P09701: Corning Tropel LightGage Metrology System
/public/

Week 6:The Build Continues

Week 6: The Build Continues (by Friday 4/17/09)

All documents generated by the P09701 team during Week 6 of MSD2 are to be posted here.

MSD2 Week 6 Progress

Tuesday's News

1. Invar post assemblies completed--set screws and vee pads loctited in place. Posts and locations labeled.

2. Hardware for mounting LG heads acquired.

3. Bad actuator being sent back to vendor to be replaced. Should have replacement late this week or early next week.

4. TMC legs have been extracted from stock room and will be transported to RIT by Friday.

5. Week 8 Imagine RIT poster nearly complete.

Thursday Progress

1. Hole was cut in the base plate, but is too small 4in accounts for flat side, to flat side of lightgage head, not the rounded sides.

None-the-less after a few hours and 3 end mills:

Image:hole1.jpg Image:hole2.jpg Image:hole3.jpg Image:hole4.jpg Image:hole5.jpg

2. Week 8 Poster Preview:

public/PosterPreview.png

Friday Progress

1. Connected LightGage to computer and verified proper operation.

2. Nick finished drilling out hole in table. Head now fits and can be rotated freely.

3. Cara and Matt worked on build documentation--build notes nearly complete and pictures were taken to accompany the text.

4. It is now possible to establish communication between all 3 actuators--addressing each one independently. Found out how to create scripts and what the commands look like. This will help a lot in incorporating the actuators into TMS.

5. Top LG head installed on top plate, and bottom "dummy" LG head was secured to the bottom plate. Using a harp, a part was placed on the bottom fizeau and using the micrometers we were able to find and null fringes on the bottom fizeau. This was much easier than expected--simply move all three micrometers to the same position and finding fringes was trivial.

Interestingly, it was found that the micrometers have a virtually zero settling time--once they were moved, there was no further movement and the part can be measured right after moving the micrometers with no ill-effects. It appears that there will be no problems with micrometer drift.

It should also be noted that there are some vibration issues that must be addressed. The TMC table has yet to be set up, so once we get it floating some of the vibration should be isolated. However, there may be issues with circulating air, as simply blowing on the top head causes the fringes to move. A worst-case scenario would mean building a makeshift enclosure around the system to prevent vibration due to air circulation. Foam board is a cost-effective and simple to implement solution to this.

6. Part was measured using top LG head. The measurement data showed clearly the different features of the part surface, including the distinct surface profiles at different heights. This is good news. However there was a little noise in the measurement--most likely due to vibration. It also seems like the laser is skipping modes, so we might not be getting the full dynamic range we should be getting. This is a question for Tropel. Matt will look at the state file of the demo LG at Tropel for insight.

That is all for week 6! See below for pictures of Friday's activities...

Friday Pictures:

Image:friday1.jpg Image:friday2.jpg Image:friday3.jpg Image:friday4.jpg Image:friday5.jpg Image:friday6.jpg Image:friday7.jpg Image:friday8.jpg Image:friday9.jpg Image:friday10.jpg Image:friday11.jpg Image:friday12.jpg Image:friday13.jpg Image:friday14.jpg Image:friday15.jpg Image:friday16.jpg Image:friday17.jpg Image:friday18.jpg Image:friday19.jpg Image:friday20.jpg Image:friday21.jpg Image:friday22.jpg